3sinA=4cosA find tan​​​ thita + cot thita/tan​​​ thita - cot thita

  1. 3sinA=4cosA find tan​​​ thita + cot thita/tan​​​ thita - cot thita

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
TanA = 4/3
And 
Cot A = ¾
 
Now
expressions becomes= 
= (3/4 + 4/3 )/(4/3 - 3/4)
= (25/12) /(7/12)
= 25/7
 
 
Regards
Arun (askIITians forum expert)